OpenStudy (anonymous):

Solve the system. The answer is (3,0)

OpenStudy (lgbasallote):

...good to know....?

OpenStudy (lgbasallote):

i assume you want to know how they got that?

OpenStudy (anonymous):

Indeed you are correct

OpenStudy (lgbasallote):

4x + 6y = 12 -10x + 9y = -30 if i multiply the first equation by 5 and the second equation by 2 i get 20x + 30y = 60 -20x + 18y = -60 if i add the two equations i have 0x + 48y = 0 if i solve for y, i get y = 0 now, if i substitute y = 0 into 4x + 6y = 12 and solve for x i get 4x - 6(0) = 12 4x = 12 x = 3 so x = 3 and y = 0 therefore the intersection is at (3,0) does that help?

OpenStudy (anonymous):

Yes, thank you.

OpenStudy (lgbasallote):

welcome